Solve. Identify double roots for the equation
(x-3)^3-(x-3)^2=0

Answer by user_dude2008(1862)

(x-3)^3-(x-3)^2=0

(x-3)^2(x-3-1)=0

(x-3)^2=0 or x-4=0

x=3 (double root) or x=4